OpenStudy (anonymous):

4 3/4 x 4 3/4 (Simplify your answer and write it as a proper fraction or as a whole or mixed number.)

jimthompson5910 (jim_thompson5910):

4 3/4 = 4+3/4 = 16/4+3/4 = 19/4 So 4 3/4 x 4 3/4 becomes 19/4 x 19/4 19/4 x 19/4 (19 x 19)/(4 x 4) 361/16 (352+9)/16 352/16+9/16 22+9/16 22 9/16 So 4 3/4 x 4 3/4 = 22 9/16

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I dont know where my head was im sorry.

OpenStudy (anonymous):

no problem i understand